OpenStudy (anonymous):

log8 base 16... i cant figure this out!

OpenStudy (anonymous):

somebody should help me please

OpenStudy (anonymous):

help please!

OpenStudy (anonymous):

\[\log_{16}8 \]

OpenStudy (anonymous):

ok think of this this way

OpenStudy (anonymous):

what this means is: 16 to what power is equal to 8?

OpenStudy (anonymous):

\[\log_{16} 8\] ---> 16^x = 8.. 16^(0.75) = 8 therefore, x = 0.75

OpenStudy (anonymous):

.75 thanks!

OpenStudy (anonymous):

16^n=8 (2^4)^n=2^3 4n=3 n=3/4

OpenStudy (anonymous):

ure welcome

OpenStudy (anonymous):

Yiano got the result first. Good work.